OpenStudy (anonymous):

find an equation of the tangent line to the graph of f at the given point. f(x)=square root of (x-1)....with points (5,2)

OpenStudy (zarkon):

\[y-y_1=f'(x_1)(x-x_1)\]

OpenStudy (anonymous):

what is that?

OpenStudy (zarkon):

the equation of the line tangent to f(x) at the point \[(x_1,y_1)\]

myininaya (myininaya):

zarkon!

OpenStudy (zarkon):

yo!

myininaya (myininaya):

i think i did something take a look at this when you are done :) http://openstudy.com/groups/mathematics/updates/4e8bae890b8b543d41faa262#/groups/mathematics/updates/4e8bb0ac0b8b543d41fad5b4

myininaya (myininaya):

i mean if you want to take a look at it

OpenStudy (zarkon):

lol

myininaya (myininaya):

you don't have to no one is gonna make you

OpenStudy (anonymous):

can you show steps cause im confused

OpenStudy (mimi_x3):

f'(x) = 1/2sqrt(x-1) m = 1/4 Tangent: y-y1 = m(x-x1) y-2 = 1/4(x-5) 4y - 8 = x -5 T: x-4y +3 = 0

OpenStudy (anonymous):

how did you get the derivative?

OpenStudy (mimi_x3):